# Do we still need Human Annotators? Prompting Large Language Models for Aspect Sentiment Quad Prediction

### v1

@@ E.2 ASQP
Aspect Term - 0 65.60 57.02 77.23 71.30 64.43 79.82 68.35 59.08 81.07 64.69 59.45 70.95 66.79 59.18 76.66
10 67.38 60.12 76.64 74.53 67.96 82.52 73.29 67.71 79.88 66.65 63.79 69.81 67.75 63.78 72.27
20 68.59 63.37 74.78 74.87 68.39 82.73 73.67 66.89 81.98 67.44 63.48 71.94 68.24 62.44 75.24
30 67.94 62.89 73.87 75.46 69.60 82.41 74.12 67.40 82.34 69.54 64.86 74.98 69.65 64.41 75.84
40 68.27 63.97 73.20 76.19 71.46 81.60 74.93 70.10 80.48 69.32 64.84 74.45 67.61 63.59 72.17
50 69.42 64.93 74.58 75.93 72.80 79.36 74.43 70.89 78.34 69.57 64.89 74.98 68.17 66.55 69.89
SC 0 67.57 62.23 73.91 74.19 71.17 77.48 68.73 63.68 74.65 66.97 64.68 69.43 68.78 63.80 74.60
10 70.68 68.39 73.12 75.94 76.34 75.53 71.63 72.80 70.50 67.33 71.05 63.98 67.81 72.89 63.39
20 69.35 70.55 68.18 77.37 76.56 78.19 75.48 76.32 74.65 69.48 72.92 66.35 69.68 74.80 65.22
30 69.24 71.19 67.39 76.62 76.83 76.42 74.08 74.30 73.86 71.27 74.29 68.48 67.72 72.09 63.84
40 70.15 72.92 67.59 77.76 79.34 76.24 74.38 77.75 71.29 71.36 74.48 68.48 66.83 74.09 60.87
50 70.53 75.45 66.21 75.66 80.85 71.10 74.97 80.32 70.30 72.07 76.05 68.48 66.14 78.37 57.21
Opinion Term - 0 62.52 60.46 64.73 65.57 64.22 66.98 51.88 53.34 50.51 57.47 58.07 56.90 44.13 46.48 42.01
10 62.29 58.89 66.10 67.53 64.50 70.87 58.66 58.71 58.60 58.60 58.13 59.08 47.95 50.52 45.64
20 56.77 55.03 58.63 68.57 66.23 71.09 58.04 58.81 57.29 58.06 57.30 58.84 50.20 51.05 49.39
30 59.05 57.00 61.26 67.86 65.79 70.08 54.89 54.80 54.98 61.24 60.46 62.05 53.09 54.07 52.15
40 59.42 56.49 62.69 67.96 66.31 69.70 57.34 57.56 57.13 59.66 58.87 60.49 51.42 52.35 50.52
50 59.74 57.19 62.53 67.93 66.48 69.46 55.79 56.27 55.31 62.21 60.84 63.63 52.02 54.29 49.94
SC 0 62.76 64.45 61.16 65.73 68.18 63.45 52.10 58.10 47.22 58.03 62.05 54.50 44.69 50.00 40.41
10 63.51 67.01 60.35 67.05 70.91 63.59 59.44 66.67 53.62 59.41 66.89 53.44 48.32 59.45 40.70
20 55.66 61.94 50.54 67.54 72.36 63.32 58.08 67.09 51.21 58.01 65.56 52.03 52.50 64.35 44.33
30 60.01 66.67 54.57 67.40 72.76 62.77 55.90 63.78 49.76 61.90 70.75 55.03 53.40 64.34 45.64
40 60.41 67.05 54.97 67.78 72.93 63.32 58.36 67.84 51.21 61.67 69.87 55.20 52.42 64.74 44.04
50 58.05 65.87 51.88 67.21 74.22 61.41 55.73 65.63 48.43 61.80 70.18 55.20 52.63 68.13 42.88
Aspect Category - 0 59.61 59.24 60.00 66.46 66.11 66.82 66.28 65.02 67.59 74.33 73.72 74.96 64.63 65.06 64.21
10 72.31 71.55 73.09 70.99 69.51 72.53 71.24 69.52 73.05 76.01 75.11 76.96 64.58 66.58 62.70
20 72.41 73.00 71.83 77.21 76.69 77.75 75.48 75.57 75.39 77.30 75.31 79.40 67.07 67.89 66.28
30 74.19 74.79 73.61 79.82 79.81 79.82 78.68 77.98 79.39 79.18 77.78 80.64 70.20 71.47 68.98
40 75.51 75.37 75.64 80.86 81.21 80.53 79.24 78.75 79.74 79.20 77.56 80.92 70.32 71.78 68.91
50 75.54 75.82 75.27 81.03 81.66 80.41 78.93 78.78 79.09 78.87 77.07 80.76 69.88 73.07 66.98
SC 0 57.86 61.55 54.58 64.35 68.03 61.05 65.33 68.99 62.04 72.63 75.70 69.80 64.96 69.11 61.27
10 71.23 78.92 64.90 69.91 75.42 65.15 70.77 77.11 65.40 72.99 81.33 66.20 63.84 75.57 55.26
20 69.93 81.37 61.32 78.05 85.79 71.60 74.92 84.71 67.15 75.00 82.14 69.00 66.41 79.35 57.10
30 73.41 85.23 64.47 79.94 88.89 72.62 77.87 87.16 70.36 76.16 84.98 69.00 67.12 80.09 57.76
40 72.98 84.82 64.04 80.64 89.32 73.50 77.90 89.10 69.20 76.81 84.78 70.20 67.85 82.89 57.43
50 73.53 88.51 62.89 79.67 90.04 71.45 77.00 88.45 68.18 76.63 85.86 69.20 66.94 84.87 55.26
Sentiment Polarity - 0 90.19 90.02 90.36 90.99 90.54 91.45 92.75 93.05 92.44 89.52 90.30 88.75 90.13 90.85 89.43
10 90.41 89.89 90.94 91.68 90.86 92.52 93.78 94.27 93.30 90.45 91.49 89.44 86.88 91.19 82.97
20 90.34 89.74 90.94 92.58 92.71 92.45 94.23 95.28 93.20 91.79 91.46 92.13 89.50 91.40 87.69
30 90.79 90.78 90.80 93.23 93.42 93.03 94.88 95.63 94.15 93.02 92.51 93.55 90.33 93.14 87.69
40 91.56 91.16 91.97 94.10 94.17 94.03 94.34 94.73 93.95 92.88 92.03 93.74 89.97 92.27 87.78
50 91.38 91.24 91.52 93.41 93.69 93.14 94.44 94.86 94.02 93.28 92.54 94.03 90.84 94.36 87.59
SC 0 86.92 91.34 82.91 88.75 92.84 85.00 87.75 94.30 82.05 87.26 92.12 82.89 87.95 92.91 83.49
10 85.66 92.64 79.66 86.62 93.95 80.34 87.92 96.33 80.85 84.64 94.29 76.77 82.38 93.96 73.35
20 83.14 93.21 75.04 87.02 94.91 80.34 87.25 97.47 78.97 85.03 93.81 77.75 81.50 94.41 71.70
30 84.04 93.88 76.07 88.01 96.31 81.03 87.09 97.06 78.97 84.71 94.85 76.53 81.63 96.46 70.75
40 83.62 94.41 75.04 87.62 96.09 80.52 85.47 96.96 76.41 86.13 94.72 78.97 81.78 95.58 71.46
50 82.01 94.43 72.48 85.96 95.57 78.10 86.82 97.45 78.29 85.79 96.06 77.51 80.72 97.02 69.10
@@
Aspect Term - 0 57.66 51.95 64.82 63.37 59.28 68.09 55.61 50.85 61.43 60.76 56.90 65.21 55.85 52.81 59.31
10 53.31 48.11 59.88 60.21 56.86 64.01 63.76 63.79 63.80 53.05 55.76 50.85 53.74 60.24 48.56
20 58.40 55.33 61.86 62.57 60.07 65.32 65.63 64.08 67.29 53.20 54.33 52.13 55.80 55.87 55.79
30 59.50 56.25 63.20 65.79 63.04 68.79 66.31 63.12 69.86 57.19 57.06 57.39 57.97 59.66 56.38
40 58.89 57.23 60.67 67.40 66.97 67.84 67.05 66.60 67.52 59.35 57.90 60.95 57.96 59.63 56.38
50 61.39 60.27 62.57 68.40 68.15 68.65 68.71 69.57 67.88 60.17 56.92 63.84 59.99 62.49 57.71
SC 0 60.56 59.13 62.06 63.87 65.79 62.06 55.24 58.08 52.67 63.56 64.10 63.03 56.57 58.58 54.69
10 48.92 77.35 35.77 48.46 87.50 33.51 55.05 81.40 41.58 43.63 82.78 29.62 42.16 80.13 28.60
20 54.64 83.06 40.71 57.99 87.19 43.44 62.83 85.37 49.70 45.81 82.21 31.75 45.89 91.16 30.66
30 57.55 84.35 43.68 62.61 85.80 49.29 62.50 84.75 49.50 56.14 85.85 41.71 50.25 88.95 35.01
40 59.49 84.67 45.85 62.73 87.34 48.94 61.08 87.45 46.93 58.54 82.05 45.50 52.65 88.17 37.53
50 59.69 86.19 45.65 63.86 88.92 49.82 62.86 87.32 49.11 59.18 82.28 46.21 54.31 89.95 38.90
Opinion Term - 0 52.53 55.87 49.57 57.27 60.43 54.43 43.26 49.94 38.16 53.28 57.82 49.42 39.94 47.49 34.48
10 50.58 48.62 52.72 54.93 50.87 59.70 47.79 48.72 46.91 48.71 49.70 47.76 37.17 43.52 32.44
20 50.39 49.31 51.53 56.71 54.44 59.18 48.58 50.44 46.86 50.90 50.20 51.64 42.90 44.81 41.16
30 50.30 48.96 51.72 59.83 58.21 61.55 47.25 48.42 46.16 55.15 54.85 55.45 43.97 47.00 41.31
40 52.35 50.60 54.25 60.10 58.73 61.55 48.45 49.56 47.42 54.79 54.97 54.64 45.45 49.28 42.18
50 54.40 53.27 55.59 60.31 58.75 61.98 49.04 50.25 47.90 57.14 56.73 57.57 47.89 52.46 44.07
SC 0 52.45 60.82 46.10 56.61 66.18 49.46 41.34 55.74 32.85 55.03 64.83 47.80 40.11 52.98 32.27
10 40.31 75.09 27.55 39.92 75.19 27.17 41.32 69.12 29.47 38.56 78.38 25.57 31.01 65.12 20.35
20 41.98 75.79 29.03 48.92 79.51 35.33 45.88 70.78 33.94 40.99 77.45 27.87 33.59 71.56 21.95
30 46.55 75.76 33.60 53.81 79.16 40.76 44.94 70.54 32.97 49.33 79.30 35.80 39.23 74.69 26.60
40 47.06 74.42 34.41 53.01 78.25 40.08 45.06 73.50 32.49 49.76 76.56 36.86 39.41 72.66 27.03
50 46.41 75.68 33.47 54.17 78.21 41.44 47.65 72.41 35.51 50.77 76.79 37.92 42.01 74.72 29.22
Aspect Category - 0 41.78 45.12 38.91 44.29 46.54 42.25 43.52 47.05 40.50 74.10 77.42 71.08 54.21 61.08 48.75
10 56.62 53.15 60.60 53.13 48.89 58.18 60.75 60.04 61.49 60.97 60.44 61.52 50.45 56.73 45.44
20 60.58 58.76 62.55 58.58 56.02 61.41 68.74 70.29 67.27 64.15 61.38 67.20 53.90 55.40 52.49
30 62.57 61.15 64.07 64.22 62.38 66.18 66.60 65.62 67.62 69.59 67.88 71.40 57.09 59.29 55.06
40 64.31 62.65 66.07 61.71 59.92 63.63 69.79 68.52 71.12 71.64 69.67 73.76 58.36 61.67 55.39
50 64.16 62.96 65.42 65.36 64.40 66.35 71.17 70.22 72.15 71.12 68.66 73.76 58.74 62.20 55.66
SC 0 39.11 45.86 34.10 40.17 46.62 35.29 38.58 47.64 32.41 73.64 82.96 66.20 50.89 62.90 42.74
10 47.45 86.69 32.66 40.89 75.00 28.11 56.35 87.93 41.46 45.60 89.47 30.60 44.28 89.80 29.38
20 50.87 89.53 35.53 49.95 81.46 36.02 61.57 89.20 47.01 50.36 89.74 35.00 45.06 93.19 29.72
30 52.39 85.95 37.68 57.14 84.97 43.05 61.27 90.08 46.42 58.34 90.72 43.00 49.70 90.71 34.22
40 57.14 90.12 41.83 54.11 79.55 41.00 61.85 93.24 46.28 61.38 90.62 46.40 49.64 87.03 34.72
50 54.89 88.25 39.83 57.88 84.31 44.07 64.28 90.69 49.78 63.01 90.94 48.20 51.94 87.70 36.89
Sentiment Polarity - 0 86.90 90.61 83.49 88.05 90.90 85.38 90.23 94.07 86.70 88.68 90.66 86.80 89.57 92.39 86.93
10 84.13 83.12 85.16 84.86 81.11 88.97 86.71 87.49 85.95 85.78 85.59 85.97 81.61 91.41 73.73
20 85.33 82.93 87.90 88.25 86.84 89.72 88.80 90.37 87.28 82.77 79.08 86.85 87.34 91.57 83.49
30 86.98 85.83 88.17 88.12 86.80 89.48 90.04 91.41 88.72 84.05 81.43 86.85 88.52 92.31 85.05
40 87.46 86.57 88.38 88.94 88.33 89.55 87.97 87.47 88.48 87.34 85.57 89.19 88.78 92.99 84.95
50 87.68 87.27 88.10 89.50 88.94 90.07 88.69 88.42 88.96 87.65 85.50 89.93 89.00 93.46 84.95
SC 0 83.29 93.04 75.38 83.35 92.99 75.52 82.75 95.53 72.99 85.45 92.09 79.71 85.23 94.54 77.59
10 56.48 97.08 39.83 53.68 97.29 37.07 64.00 96.55 47.86 53.52 95.60 37.16 56.86 97.70 40.09
20 59.88 97.31 43.25 61.81 97.05 45.34 71.30 97.91 56.07 59.08 96.67 42.54 55.16 97.60 38.44
30 61.95 96.39 45.64 67.19 96.45 51.55 68.94 96.90 53.50 64.72 95.69 48.90 61.61 97.45 45.05
40 65.15 97.61 48.89 65.46 94.77 50.00 68.74 98.41 52.82 68.13 95.18 53.06 64.56 98.08 48.11
50 63.22 96.49 47.01 68.88 96.28 53.62 72.47 97.68 57.61 70.09 96.57 55.01 67.39 97.76 51.42

### v2

@@ E.2 ASQP
Aspect Term - 0 65.60 57.02 77.23 71.30 64.43 79.82 64.69 59.45 70.95 54.57 43.54 73.06 66.79 59.18 76.66
10 67.38 60.12 76.64 74.53 67.96 82.52 66.65 63.79 69.81 63.83 54.41 77.18 67.75 63.78 72.27
20 68.59 63.37 74.78 74.87 68.39 82.73 67.44 63.48 71.94 67.31 58.83 78.69 68.24 62.44 75.24
30 67.94 62.89 73.87 75.46 69.60 82.41 69.54 64.86 74.98 68.21 59.23 80.41 69.65 64.41 75.84
40 68.27 63.97 73.20 76.19 71.46 81.60 69.32 64.84 74.45 68.19 60.18 78.69 67.61 63.59 72.17
50 69.42 64.93 74.58 75.93 72.80 79.36 69.57 64.89 74.98 70.31 63.17 79.31 68.17 66.55 69.89
SC 0 67.57 62.23 73.91 74.19 71.17 77.48 66.97 64.68 69.43 59.05 51.96 68.38 68.78 63.80 74.60
10 70.68 68.39 73.12 75.94 76.34 75.53 67.33 71.05 63.98 62.25 64.93 59.79 67.81 72.89 63.39
20 69.35 70.55 68.18 77.37 76.56 78.19 69.48 72.92 66.35 66.42 72.18 61.51 69.68 74.80 65.22
30 69.24 71.19 67.39 76.62 76.83 76.42 71.27 74.29 68.48 65.91 73.42 59.79 67.72 72.09 63.84
40 70.15 72.92 67.59 77.76 79.34 76.24 71.36 74.48 68.48 65.23 76.15 57.04 66.83 74.09 60.87
50 70.53 75.45 66.21 75.66 80.85 71.10 72.07 76.05 68.48 67.98 79.36 59.45 66.14 78.37 57.21
Opinion Term - 0 62.52 60.46 64.73 65.57 64.22 66.98 57.47 58.07 56.90 29.95 26.72 34.07 44.13 46.48 42.01
10 62.29 58.89 66.10 67.53 64.50 70.87 58.60 58.13 59.08 41.73 38.03 46.24 47.95 50.52 45.64
20 56.77 55.03 58.63 68.57 66.23 71.09 58.06 57.30 58.84 43.16 39.66 47.33 50.20 51.05 49.39
30 59.05 57.00 61.26 67.86 65.79 70.08 61.24 60.46 62.05 44.18 40.78 48.19 53.09 54.07 52.15
40 59.42 56.49 62.69 67.96 66.31 69.70 59.66 58.87 60.49 43.79 40.65 47.47 51.42 52.35 50.52
50 59.74 57.19 62.53 67.93 66.48 69.46 62.21 60.84 63.63 43.10 39.80 47.01 52.02 54.29 49.94
SC 0 62.76 64.45 61.16 65.73 68.18 63.45 58.03 62.05 54.50 29.09 28.99 29.19 44.69 50.00 40.41
10 63.51 67.01 60.35 67.05 70.91 63.59 59.41 66.89 53.44 41.79 49.84 35.97 48.32 59.45 40.70
20 55.66 61.94 50.54 67.54 72.36 63.32 58.01 65.56 52.03 43.05 52.61 36.43 52.50 64.35 44.33
30 60.01 66.67 54.57 67.40 72.76 62.77 61.90 70.75 55.03 43.57 54.21 36.43 53.40 64.34 45.64
40 60.41 67.05 54.97 67.78 72.93 63.32 61.67 69.87 55.20 40.23 53.16 32.35 52.42 64.74 44.04
50 58.05 65.87 51.88 67.21 74.22 61.41 61.80 70.18 55.20 42.92 54.74 35.29 52.63 68.13 42.88
Aspect Category - 0 59.61 59.24 60.00 66.46 66.11 66.82 74.33 73.72 74.96 42.14 39.67 44.96 64.63 65.06 64.21
10 72.31 71.55 73.09 70.99 69.51 72.53 76.01 75.11 76.96 40.05 38.39 41.86 64.58 66.58 62.70
20 72.41 73.00 71.83 77.21 76.69 77.75 77.30 75.31 79.40 44.53 43.17 45.99 67.07 67.89 66.28
30 74.19 74.79 73.61 79.82 79.81 79.82 79.18 77.78 80.64 47.28 46.30 48.31 70.20 71.47 68.98
40 75.51 75.37 75.64 80.86 81.21 80.53 79.20 77.56 80.92 47.06 46.03 48.14 70.32 71.78 68.91
50 75.54 75.82 75.27 81.03 81.66 80.41 78.87 77.07 80.76 49.09 48.30 49.92 69.88 73.07 66.98
SC 0 57.86 61.55 54.58 64.35 68.03 61.05 72.63 75.70 69.80 43.33 46.05 40.91 64.96 69.11 61.27
10 71.23 78.92 64.90 69.91 75.42 65.15 72.99 81.33 66.20 41.45 52.37 34.30 63.84 75.57 55.26
20 69.93 81.37 61.32 78.05 85.79 71.60 75.00 82.14 69.00 45.04 58.61 36.57 66.41 79.35 57.10
30 73.41 85.23 64.47 79.94 88.89 72.62 76.16 84.98 69.00 47.47 63.76 37.81 67.12 80.09 57.76
40 72.98 84.82 64.04 80.64 89.32 73.50 76.81 84.78 70.20 42.44 59.26 33.06 67.85 82.89 57.43
50 73.53 88.51 62.89 79.67 90.04 71.45 76.63 85.86 69.20 46.56 64.71 36.36 66.94 84.87 55.26
Sentiment Polarity - 0 90.19 90.02 90.36 90.99 90.54 91.45 89.52 90.30 88.75 88.09 87.60 88.59 90.13 90.85 89.43
10 90.41 89.89 90.94 91.68 90.86 92.52 90.45 91.49 89.44 87.04 87.07 87.02 86.88 91.19 82.97
20 90.34 89.74 90.94 92.58 92.71 92.45 91.79 91.46 92.13 88.84 89.00 88.68 89.50 91.40 87.69
30 90.79 90.78 90.80 93.23 93.42 93.03 93.02 92.51 93.55 88.99 89.07 88.93 90.33 93.14 87.69
40 91.56 91.16 91.97 94.10 94.17 94.03 92.88 92.03 93.74 89.46 89.11 89.80 89.97 92.27 87.78
50 91.38 91.24 91.52 93.41 93.69 93.14 93.28 92.54 94.03 89.34 89.36 89.32 90.84 94.36 87.59
SC 0 86.92 91.34 82.91 88.75 92.84 85.00 87.26 92.12 82.89 81.15 87.57 75.61 87.95 92.91 83.49
10 85.66 92.64 79.66 86.62 93.95 80.34 84.64 94.29 76.77 72.14 90.44 60.00 82.38 93.96 73.35
20 83.14 93.21 75.04 87.02 94.91 80.34 85.03 93.81 77.75 74.45 94.38 61.46 81.50 94.41 71.70
30 84.04 93.88 76.07 88.01 96.31 81.03 84.71 94.85 76.53 70.87 92.19 57.56 81.63 96.46 70.75
40 83.62 94.41 75.04 87.62 96.09 80.52 86.13 94.72 78.97 70.73 94.31 56.59 81.78 95.58 71.46
50 82.01 94.43 72.48 85.96 95.57 78.10 85.79 96.06 77.51 70.52 93.55 56.59 80.72 97.02 69.10
@@
Aspect Term - 0 57.66 51.95 64.82 63.37 59.28 68.09 60.76 56.90 65.21 38.49 31.52 49.42 55.85 52.81 59.31
10 53.31 48.11 59.88 60.21 56.86 64.01 53.05 55.76 50.85 44.81 40.58 50.17 53.74 60.24 48.56
20 58.40 55.33 61.86 62.57 60.07 65.32 53.20 54.33 52.13 53.08 50.41 56.29 55.80 55.87 55.79
30 59.50 56.25 63.20 65.79 63.04 68.79 57.19 57.06 57.39 56.07 53.93 58.63 57.97 59.66 56.38
40 58.89 57.23 60.67 67.40 66.97 67.84 59.35 57.90 60.95 60.19 58.85 61.65 57.96 59.63 56.38
50 61.39 60.27 62.57 68.40 68.15 68.65 60.17 56.92 63.84 58.04 56.22 60.00 59.99 62.49 57.71
SC 0 60.56 59.13 62.06 63.87 65.79 62.06 63.56 64.10 63.03 41.48 38.97 44.33 56.57 58.58 54.69
10 48.92 77.35 35.77 48.46 87.50 33.51 43.63 82.78 29.62 15.67 89.29 8.59 42.16 80.13 28.60
20 54.64 83.06 40.71 57.99 87.19 43.44 45.81 82.21 31.75 29.63 86.67 17.87 45.89 91.16 30.66
30 57.55 84.35 43.68 62.61 85.80 49.29 56.14 85.85 41.71 40.32 92.59 25.77 50.25 88.95 35.01
40 59.49 84.67 45.85 62.73 87.34 48.94 58.54 82.05 45.50 38.02 95.83 23.71 52.65 88.17 37.53
50 59.69 86.19 45.65 63.86 88.92 49.82 59.18 82.28 46.21 40.21 91.46 25.77 54.31 89.95 38.90
Opinion Term - 0 52.53 55.87 49.57 57.27 60.43 54.43 53.28 57.82 49.42 29.21 28.16 30.36 39.94 47.49 34.48
10 50.58 48.62 52.72 54.93 50.87 59.70 48.71 49.70 47.76 29.17 26.09 33.08 37.17 43.52 32.44
20 50.39 49.31 51.53 56.71 54.44 59.18 50.90 50.20 51.64 31.01 28.42 34.12 42.90 44.81 41.16
30 50.30 48.96 51.72 59.83 58.21 61.55 55.15 54.85 55.45 33.63 31.13 36.56 43.97 47.00 41.31
40 52.35 50.60 54.25 60.10 58.73 61.55 54.79 54.97 54.64 36.18 34.20 38.42 45.45 49.28 42.18
50 54.40 53.27 55.59 60.31 58.75 61.98 57.14 56.73 57.57 35.96 33.60 38.69 47.89 52.46 44.07
SC 0 52.45 60.82 46.10 56.61 66.18 49.46 55.03 64.83 47.80 29.92 34.41 26.47 40.11 52.98 32.27
10 40.31 75.09 27.55 39.92 75.19 27.17 38.56 78.38 25.57 12.35 68.18 6.79 31.01 65.12 20.35
20 41.98 75.79 29.03 48.92 79.51 35.33 40.99 77.45 27.87 17.19 62.86 9.95 33.59 71.56 21.95
30 46.55 75.76 33.60 53.81 79.16 40.76 49.33 79.30 35.80 26.74 70.19 16.52 39.23 74.69 26.60
40 47.06 74.42 34.41 53.01 78.25 40.08 49.76 76.56 36.86 26.82 75.79 16.29 39.41 72.66 27.03
50 46.41 75.68 33.47 54.17 78.21 41.44 50.77 76.79 37.92 27.59 69.72 17.19 42.01 74.72 29.22
Aspect Category - 0 41.78 45.12 38.91 44.29 46.54 42.25 74.10 77.42 71.08 44.76 46.86 42.85 54.21 61.08 48.75
10 56.62 53.15 60.60 53.13 48.89 58.18 60.97 60.44 61.52 25.64 23.92 27.64 50.45 56.73 45.44
20 60.58 58.76 62.55 58.58 56.02 61.41 64.15 61.38 67.20 32.97 31.64 34.42 53.90 55.40 52.49
30 62.57 61.15 64.07 64.22 62.38 66.18 69.59 67.88 71.40 33.27 31.85 34.83 57.09 59.29 55.06
40 64.31 62.65 66.07 61.71 59.92 63.63 71.64 69.67 73.76 34.74 34.11 35.41 58.36 61.67 55.39
50 64.16 62.96 65.42 65.36 64.40 66.35 71.12 68.66 73.76 34.67 33.70 35.70 58.74 62.20 55.66
SC 0 39.11 45.86 34.10 40.17 46.62 35.29 73.64 82.96 66.20 42.22 52.45 35.33 50.89 62.90 42.74
10 47.45 86.69 32.66 40.89 75.00 28.11 45.60 89.47 30.60 7.18 42.22 3.93 44.28 89.80 29.38
20 50.87 89.53 35.53 49.95 81.46 36.02 50.36 89.74 35.00 16.22 63.38 9.30 45.06 93.19 29.72
30 52.39 85.95 37.68 57.14 84.97 43.05 58.34 90.72 43.00 24.79 69.52 15.08 49.70 90.71 34.22
40 57.14 90.12 41.83 54.11 79.55 41.00 61.38 90.62 46.40 22.45 68.42 13.43 49.64 87.03 34.72
50 54.89 88.25 39.83 57.88 84.31 44.07 63.01 90.94 48.20 24.58 66.36 15.08 51.94 87.70 36.89
Sentiment Polarity - 0 86.90 90.61 83.49 88.05 90.90 85.38 88.68 90.66 86.80 84.41 87.53 81.51 89.57 92.39 86.93
10 84.13 83.12 85.16 84.86 81.11 88.97 85.78 85.59 85.97 85.59 82.76 88.63 81.61 91.41 73.73
20 85.33 82.93 87.90 88.25 86.84 89.72 82.77 79.08 86.85 89.29 88.73 89.85 87.34 91.57 83.49
30 86.98 85.83 88.17 88.12 86.80 89.48 84.05 81.43 86.85 86.41 85.07 87.80 88.52 92.31 85.05
40 87.46 86.57 88.38 88.94 88.33 89.55 87.34 85.57 89.19 88.30 88.03 88.59 88.78 92.99 84.95
50 87.68 87.27 88.10 89.50 88.94 90.07 87.65 85.50 89.93 88.89 88.81 88.98 89.00 93.46 84.95
SC 0 83.29 93.04 75.38 83.35 92.99 75.52 85.45 92.09 79.71 77.82 90.88 68.05 85.23 94.54 77.59
10 56.48 97.08 39.83 53.68 97.29 37.07 53.52 95.60 37.16 17.66 93.02 9.76 56.86 97.70 40.09
20 59.88 97.31 43.25 61.81 97.05 45.34 59.08 96.67 42.54 28.81 100.00 16.83 55.16 97.60 38.44
30 61.95 96.39 45.64 67.19 96.45 51.55 64.72 95.69 48.90 37.25 95.00 23.17 61.61 97.45 45.05
40 65.15 97.61 48.89 65.46 94.77 50.00 68.13 95.18 53.06 36.90 98.94 22.68 64.56 98.08 48.11
50 63.22 96.49 47.01 68.88 96.28 53.62 70.09 96.57 55.01 40.54 97.22 25.61 67.39 97.76 51.42